LONG TITLE

NAME SURNAME

Abstract. In this environment a summary should be put.

If the Polish characters do not appear the following commands might be used: “k–a˝ “k–e˝ “’–c˝ “’–n˝ “’–o˝ “’–s˝ “’–z˝ “.–z˝ –“l˝ for the letters ą ę ć ń ó ś ź ż ł.

1. Available environments

1.1. Theorem type. Theorem 1 (theorem’s author, year). Theorem’s content should be written in the theorem environment.

To make reference marks for theorems, lemmas and so on you should create labels with a command “label–label’s˙ name˝ . This is the way to avoid mistakes in numbering theorems in the other parts of a lecture.

To put a reference mark for a theorem you should write “ref–label’s˙ name˝ . Here we refer to Theorem 1. Lemma 2. This is the environment for lemmas (lemma). Sublemma 3. You may have sublemmas. They should be written in the sublemma environment. Statement 4. The statement environment is intended for statements. Fact 5. The fact environment is for facts. Proposition 6. In the proposition environment you should write propositions. Corollary 7. The corollary environment should contain corollaries. Corollary. The (corollary*) environment is for unnumbered corollaries.

A similar rule applies to theorems and propositions. Hypothesis 8 (author, year). The hypothesis environment is for hypotheses.

You can write that the hipothesis is taken from a book mentioned in the bibliography in the place [2]. A command “cite–number˙ in˙ the˙ bibliography˝ may be helpful then. Observation 9. For observations there is the observation environment. Property 10. The property environment is for properties. Exercise 11. In this place exercises for readers can be left (exercise).

1.2. Definition type. The second environment type has a different style and should be used for:

Definition 12. The (definition) environment. The numeration is continued although we have different type of the environment.

It is good to emphasize a defined object with italic.

Example 13. A similar style has the (example) environment which should be used for examples.

Example. Unnumbered examples should be written in (example*).

1.3. Other.

Remark 14. There is also the environment type for remarks (remark).

Proof. The proof environment makes the text clear. Proofs are ended with a square on the left-hand side. □

2. Images

2.1. Including external images. All images must be black and white, moreover they should be in a .eps file. We suggest that coloured lines in the graphs are replaced e.g. with dashed, dotted or bold ones.

Images should be put in the figure environment. To insert an image you can use a command “includegraphics:

“begin–figure˝ [ht]

“includegraphics[height=0.5“textheight]–Image1.eps˝

“caption–Example of labeling an image˝

“end–figure˝

Figure 1. Example of labeling an image

You can change the image size by increasing or decreasing number in the square brackets. It is good to scale the image in proportion to the text size. You can scale in respect to a text width or hight:

• width=fraction“textwidth

• height=fraction“textheight

2.2. Creating images in a .tex file. Images created in TEX are of course also possible. Please remember to use only two colours: black and white.

[figure]

Figure 2. Example of labeling an image

number example
1  abc
2  def

Table 1. Example of a table

2.3. Tables. The table environment enables you to insert a table. You can refer to the table, label it or give it a number.

References

[1] N.Surname, Title, where the article was published, pages, publishing company, year.

[2] N2.Surname2, Title2, where the article was published, pages, publishing company, year.

Name Surname, (Faculty, Institute) University, street, post code town (country)

E-mail address: login@domain.rest.com